FEL Upgrade Project Weekly Brief
April 9-13, 2001

We made good progress this week on the design of the optical cavity and various magnet systems.

A requested summary of our presentation to the Defense Science Board High Energy Laser Panel on Jan. 25th was forwarded to the DSB support office this week.  We started preparation of the March financial reports with the receipt of the project accounting reports. At the midpoint of the fiscal year with six months remaining on the Phase 1 project we are looking carefully at the spending curves to ensure a smooth landing on Oct. 1.

On Friday we were visited by Dr. Al  Skolnick for a tour a and briefing on the FEL program. Dr. Skolnick was involved with the management  of the Navy laser development program in the 1970-80's.

George Neil and Michelle Shinn were involved in numerous discussions this week with Joint Technology Office (JTO) in the preparation of the final FEL-materials exposure plan for the JTO study funded for this year.

WBS 4 (Injector):
A review of the light box layout, revealed a correction allowing for a 2" aperture in the light box by going to a 4" valve.  Drawings were completed for the cathode shield. We will build and test the door in UHV conditions.

WBS 5 (SRF):
Helium Vessels received from PHPK are being qualified this and next week.  All items leak check so far are leak tight.

WBS 8 (I&C):
We are still working on ordering parts for the ion pump power supplies.  Deborah completed re-certification for Adult CPR and AED operator. The new circuit boards for the OPBMs are out for prototype manufacturing.
Laser Safety System trouble shooting has taken a lot of time this week. Thanks to everyone for their help.
The desoldering station in lab 5 is up and running. Mike has been working on the LAN and helping with cabinet drawings. He and Dan completed their laser safety training. Everyone has been trained on how to conduct a sweep of the tunnel as well.

EECAD drawings for the Charge Totalizer Current Monitor and the LSS Alignment Mode Protection Card were signed off as complete.  The Charge Dump Current Monitor Chassis Fabrication, Assembly and Wiring Diagrams are out to be checked.  The picomotor controller chassis assembly and fabrication drawings are also complete.

WBS 9 (Transport):
Optical Chicane Dipole (DW)
o We are working on the steel procurement package
Arc Dipoles (GY, GX, GQ)
o  AES continues progress in designing the 180 degree dipole (GY). They completed design of the path length
    corrector’s connections, hoses, cover and manifolds.
o  The magnetic analysis of the GY with path length correctors ran into a snag as Dave Douglas found that
    increasing the bulk field to correct for the lower field in the path length corrector slot does not meet the field
    tolerance of 1 part in 1000 imposed by dispersed beam.  At weeks end Tom Schultheiss is calculating how
    much face steel is needed and how much parallel offset there is to the beam if it is brought into the magnet
    perpendicular but off set and comes out in a similar condition.  David says he can live with a small amount
    (say 1.5 mm) of off set.  Preliminary results however yield 3.7 mm of off set.  This problem may still not be
o  The next thing to model in the GY is two sets of holes through the yoke plates, near the coils at around
    45 degrees. These holes will accommodate two periscopes that look at each other through windows in the
    outside surfaces of the vacuum chamber.  The two periscopes will allow a HeNe Laser to establish known
    optical path that coincides with the synchrotron light generated at the apex of the magnet when the beam is
    centered.  The holes through the yoke plates are similar to the single set used in the IR Demo.  That set
    produced no measurable magnetic anomaly.
QG (3 inch quad)
The coils manufacturing specification was signed sent to Procurement.
QX (Trim Quad)
o  We continued to debug our magnetic model.
o  We will purchase conductor with a known diameter to be incorporated into the drawings and the design.
    This is necessary because of build up of tolerance in this multi turn conductor geometry dominated magnet
    will ruin the field gradient quality.
o  We continued laying out the girders and stands
o  We designed more of the beam chamber in the arcs, firming up the parameters associated with chamber

WBS 10 (Wiggler):
The dispersion section has been removed from its base to allow for the drilling of some holes the manufacturer left out.  Analysis of the measurements so far continues.

WBS 11 (Optics):
The optical cavity assembly conceptual design was reviewed and discussed; several good points were raised regarding internally supporting the linear translation, and the stiffness of the mirror mounts.  These comments are being incorporated in the design. The ultraviewer (alignment) design for 4" tubing is being detailed. We also reviewed it and came up with an action item to look at an alternate baseplate fabrications scheme.  We agreed on a particular mirror orientation (out of two possible).  We also agreed to add an extra cross (with no components) next to the ultraviewers, to allow for other optical components/diagnostics to be added at a future date.  Top-level specs on the diagnostic beam dump for the FEL were distributed, and the requisite make vs buy comparison is being made.  Procurement awarded the contract to build two prototype deformable mirrors.

Other parts for the manually-actuated safety shutter arrived and were fitted, the design works well.  A few modifications to hold the limit switches need to be made, and then we can try it in User Lab 3.  Design for the manually-actuated turning mirror progresses.  The LSOP for User Lab 2 was submitted for sign off.


No operations this week.  Planning continues for the up-coming runs to support the JTO proposal and a call for proposals for the summer run is being prepared.  We heard from Peter Schaaf (Univ. of Goettingen) that the metal nitrides samples that he made during the recent winter run appear to be very high quality nitrides.